y2 dx — x(2x + 3y)dy = 0


1
Expert's answer
2021-05-07T10:17:49-0400

We have given the differential equation,

"y^2dx-x(2x+3y)dy = 0"


"y^2dx = x(2x+3y)dy"


"\dfrac{dy}{dx} = \dfrac{y^2}{x(2x+3y)}"


Putting "y = vx"


Hence, "\dfrac{dy}{dx} = v+ x \dfrac{dv}{dx}"


Therefore,


"v+ x \dfrac{dv}{dx} = \dfrac{v^2x^2}{x(2x+3vx)}"


"x \dfrac{dv}{dx} = \dfrac{-2v^2-2v}{3v+2}"


"x \dfrac{dv}{dx} = \dfrac{-v(2v+2)}{3v+2}"


"-\int \dfrac{(3v+2)dv}{2v(v+1)} = \int \dfrac{dx}{x}"


"-\dfrac{ln(v+1)}{2} - lnv = lnx + lnC"


Replacing "v = \dfrac{y}{x}"

We get,


"-\dfrac{ln(\dfrac{y}{x}+1)}{2} - ln\dfrac{y}{x} = lnx + lnC"


This is the required solution of given differential equation.